solve by any method
9x=-22+7y
-4x+2y=2

Answer by Boreal(15235)

9x=-22+7y
-4x+2y=2
--------------
9x-7y= -22
-4x+2y=2
------------
18x-14y= -44
-28x+14y=14; multiplying the top by 2 and the bottom by 7.
-10x=-30, adding the two.
x=3
substitute into either, I will use the second
-12+2y=2
2y=14
y=7
(3,7)
Check using first
27=-22+49
x=3